# Signature Kernel Scoring Rule as Spatio-Temporal Diagnostic for Probabilistic Forecasting

### v1

@@ 2.4 Signature Augmentations
As previously mentioned, the current framing of the signature uniquely characterises a path up to translation and time parametrisation. Both of these pose issues for use as an evaluation metric and for use in a scoring rule. Learning the behaviour of weather path curvature is not enough if the forecast has significant bias via translation shifts. Weather has clear temporal cycles on the level of days, months, and years. The signature needs to relate path behaviour to time, to evaluate a forecast. We introduce the following augmentations (Morrill et al., 2021). The basepoint augmentation $\phi^{b}:\mathcal{S}(\mathbb{R}^{d})\mapsto\mathcal{S}(\mathbb{R}^{d})$ is defined as:
@@
This augmentation adds another dimension to the path, which carries the time of the observation at that point. This allows for different or irregular sampling between forecast and observations. This is critical for weather forecasting, as forecasts and observations are regularly predicted and sampled at different intervals. Moreover, the effect of missing values is mitigated, as the time dimension aligns true positions with all existing values correctly. These augmentations combined remove any invariances and guarantee the uniqueness of the signature. This property is critical for the performance of a scoring rule involving the signature kernel, as it ensures strict propriety. This will be further discussed in the following subsection.
As the signature extracts information about enclosed areas with the path, this formation directly aligns with this ability by causing loops to appear from the data. This representation benefits the capturing of certain statistics, particularly the quadratic variation, which can be represented as signature terms of the lead-lagged path (Chevyrev and Kormilitzin, 2016):
[…]
An additional choice of windowing operation can be taken. The global window is considered naturally, but changing window size via sliding, expanding or hierarchical dyadic windows changes how data is read to capture information at different scales. There is evidence that the hierarchical dyadic windowing outperforms the global window. However, during implementation we observed that large path lengths posed challenges for any strategy of weather forecasting. Over the smaller path lengths that were considered, the global window was deemed suitable. Further investigation into this hyperparameter choice would be a valid extension.
$f(\mathbf{x})-L(S(\mathbf{x}))||\leq\epsilon$
The standard signature provides a natural basis for functions of the time series. In practice, there is no consensus on which version is most suited for a machine learning task. Additionally, as will be shown in the following section, we do not need to be concerned with lowering the feature dimension of the signature, and so we use the standard signature.
@@ 2.5 The Signature Kernel Scoring Rule
The previous sections introduced the idea, intuition, and practical concerns with the signature of a path. Nonetheless, the signature remains an infinite sequence of real numbers, which still can’t be used in its current form. Practically using this signature often involves truncating at a specific order depth, where the number of terms for a $d$-dimensional path in a $k$-order truncated signature is equal to:
$\sum_{m=0}^{k}d^{m}=\frac{d(d^{k}-1)}{d-1}$
However, truncation necessarily loses some information. The exponential growth of the signature enforces harsh practical constraints on the truncation level. Given a modest path of 10 dimensions, considering the signature to a depth of only 6 is already over a million terms. There is no guarantee that computed levels contain sufficient information, with the uniqueness of the signature no longer ensured by augmentations. This leads to the signature kernel, to practically incorporate all of the information contained within the signature of a path.
The primary concept involved is the kernel trick, which allows data to be analysed in a higher (or infinitely high) dimensional space, without explicitly computing the mapping, i.e. the full signature. A kernel is a positive semidefinite function that takes vectors in the original spaces as inputs, and returns the inner product of the vectors in the feature space (Wilimitis, 2019). We can define a mapping $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ to be a kernel if there exists a Hilbert space $\mathcal{H}$ (feature space), and a feature map $\phi:\mathcal{X}\rightarrow\mathcal{H}$, such that for all $x,z\in\mathcal{X}$ we have:
@@
which maps to a single score. However, to ensure that our uplifted space is a reproducing kernel Hilbert space (RKHS), we use a static kernel on our paths to lift themselves from any topological space $\mathcal{Y}$ to a Hilbert space $\mathcal{H}$ (Xu et al., 2023). It’s crucial to be in the Hilbert space to have a complete inner product structure that guarantees the kernel trick to be applicable. Moreover, as sequential information often takes values in non-trivial ambient spaces, lifting the underlying space would be a good learning strategy regardless before considering the signature (Salvi et al., 2021). Given an RKHS($\mathcal{H}$, $k$) with a kernel $k$ on $\mathcal{Y}$, the reproducing property lets us define a general static kernel as follows:
@@
In practice, this kernel is the solution of a Goursat partial differential equation (Salvi et al., 2021). In fact, the previous condition of piecewise linearity was a necessary condition to be solved by a Goursat PDE. While other methods of solving the kernel exist, involving sequentialised kernels, the time complexity is at least $\mathcal{O}(dl^{2})$, for a path with dimension $d$ and length $l$ (Kiraly and Oberhauser, 2019). Provided the number of GPU threads exceeds the size of the discretisation grid chosen for PDE evaluation on (controlled by dyadic order), the time complexity reduces to $\mathcal{O}(dl)$ (Salvi et al., 2021). This breaks the quadratic relationship with path length, making the Goursat PDE approach the most suitable method for implementation, used by the associated sigkernel python package. Finally, we can define the signature kernel score, considering path data over the time $1:t$, and across longitude and latitude $i,j$ as follows:
$\phi(F_{i,j,1:t},o_{i,j,1:t})=E_{X,X^{\prime}\sim F_{i,j,1:t}}\left[K^{Sig}(X,X^{\prime})\right]-2E_{X\sim F_{i,j,1:t}}\left[K^{Sig}(X,o_{i,j,1:t})\right]$ (4)
Using this score, we are considering both the spread of the forecasted paths in the first term and the closeness, measured by the signature kernel, of the forecasted paths to the true observed paths in the second term. We can note the similarity between this kernel score and the energy score, with the apparent sign switch as a result of kernels measuring similarity versus distance norms representing dissimilarity. Via uniqueness guaranteed by the signature augmentations, the signature kernel score exhibits the desired fundamental properties of scoring rules, particularly being strictly proper, which ensures truthful predictions (Issa et al., 2023). Moreover, the signature kernel has been shown to be highly effective in other domains, with improved accuracy in classification on a variety of multivariate time series datasets (Salvi et al., 2021).
Additionally, we will define the kernel distance as the primary metric to be used on deterministic forecasts. Using the fact that the kernel defines an inner product distance on X and Y within the transformed space $\langle S(k_{x}),S(k_{y})\rangle$, we can instead consider the Euclidean distance on this space (Phillips and Venkatasubramanian, 2011).
[…]
Simplifying:
$d^{2}(x,y)=K^{Sig}(X,X)+K^{Sig}(Y,Y)-2K^{Sig}(X,Y)$
Without a probabilistic ensemble forecast, the distance is without expectation, with the distance across time $1:t$ and longitude, latitude $i,j$ as:
@@
We now have a distance measure, as opposed to a kind of similarity measure calculated using the inner product of $\langle S(k_{x}),S(k_{y})\rangle$. Under this measure, all values are non-negative, with the distance equalling zero when observations and forecasts have the same path. Ignoring expectations, the difference between the score and distance is the addition of the term $K^{Sig}(Y,Y)$ which represents the self-similarity of the observation path. As will be shown, this will improve the interpretability of deterministic forecasts by generally making distance monotonic, with a greater focus on forecasting skill than forecast spread.
@@
for all $\mathbf{x},\mathbf{x^{\prime}}\in\mathbb{R}^{d}$. The RBF kernel changes how the data points in the sequence are related, by introducing a notion of locality, scaled by the hyperparameter $\sigma$. This changes how close values are in kernel space, which results in different behaviours being captured by the signature kernel. The RBF may be better suited to the non-linear relationships present in weather data, but requires greater consideration during implementation to correctly tune $\sigma$. When using the RBF, $\sigma$ was chosen to be equal to 1, following positive experimental results (Issa et al., 2023). Other kernels can be used, such as the more interpretable linear kernel, but it suffers significantly from numerical instability and requires much greater attention to the scale of the input data.
Moreover, dyadic order for the PDE solver needs to be chosen. Higher dyadic orders result in higher accuracy, at the cost of calculation speed. Given our inputs are significantly scaled, such that a maximum value is roughly equal to 1, experimental results show that dyadic orders of 0 or 1, are likely sufficient (Salvi et al., 2021). We implement a dyadic order of 1, following previous work (Issa et al., 2023).
The use of the Goursat PDE solver approach also affects what weather variables can be considered. In the setting of piecewise-linear fixed dimensional paths, which move in all dimensions, typically none of the elements of the signature are zero. This is referred to as a dense signature. On the other hand, rainfall data is generally zero-inflated, particularly for certain ranges of latitude-longitude coordinates. As a result, this may result in a sparse signature, with many terms equal to zero. Methods which handle sparse signatures require significant overhead, which can be avoided if we stay in contexts resulting in dense signatures (Reizenstein and Graham, 2018).
The signature kernel scoring rule has now been fully introduced. We now move to an exploration of the properties enjoyed by this new scoring rule, demonstrated through evaluation in weather scorecards.
@@
The signature kernel scoring rule and distance on probabilistic and deterministic forecasts respectively are implemented in a notably different way than standard scores. On WeatherBench 2, scores such as RMSE are calculated at every prediction time lag. Most models provide 30 to 60 values of forecasted data from their initialisation time, spaced by twelve or six hours apart, equivalent to 360 hours or 15 days. The maximal forecast lead time for a model corresponds to when performance is no longer significantly different from predicting the average climatology. These values are averaged across all initialisation times and all locations using latitude weighting, to report score against time lag. The ECMWF then calculates their headline scores by considering the maximal forecast lead time required to reach a certain score (ECMWF, 2022).
@@
The signature kernel score can not be calculated this way. A path is a temporally related sequence, not a collection of values across space at one particular time. To better use the properties of the signature, the signature kernel score is instead calculated across chronologically sequential paths of length $k$ of the predicted data. Instead of having a value calculated for each prediction time lag, a value is determined for each path length from the initialisation time. Similarly to the previous scores, these values can be averaged across all initialisation times. Further details of implementation for model evaluation use can be found in Appendix B.
@@
The following two scorecards are created from a 10% initialisation sample evenly distributed across the year 2020. Scorecard variables are chosen based on the variable overlap between models of the ECMWF headline variables (ECMWF, 2024b). In Figure 2 we compare Fuxi and IFS ENS Mean across deterministic metrics and signature kernel distance with 60 timesteps of six hours, equivalent to 15 days. In Figure 3 we compare Neural GCM and IFS ENS across probabilistic metrics and signature kernel score with 30 timesteps of twelve hours, equivalent to 15 days. In both cases, IFS is taken as the baseline, with normalised percentage differences in metrics against its comparison being shown in colour. Darker blues signify greater positive performance for the target model as opposed to the baseline, with the opposite true of darker reds. Following ECMWF standards, calculations are split by region, with the Northern Hemisphere containing the latitude slices from 20 to 90 degrees, Tropics containing -20 to 20, and the Southern Hemisphere containing -90 to -20. There is no overlap between regions.
@@
In figure 2, the domain for ACC and MSE are the 60 future lead times, while signature kernel scoring rule (SIGK) is displayed against path length between 2 and 59. Despite this difference, there is general agreement across metrics over model performance at difference times. However, some notable behavioural differences are detected. The signature kernel distance measure has generally sharper discrimination for positive Fuxi performance, seen clearly in the Southern Hemisphere evaluation. The signature kernel also changes model performance evaluations in certain areas, unanticpated by the standard metrics. Northern and Southern Hemisphere 10 meter wind speed improves for FUXI performance for longer path lengths with signature kernel evaluation, differing from worse performance at later time leads for ACC and MSE. This implies model predictions have resulted in more appropriate path behaviours, despite specific values being worse.
@@
In Figure 3, the domain for CRPS and ENS MSE are the 30 future lead times, while SIGK is displayed against path length between 2 and 29. In this scorecard, Neural GCM demonstrates superior performance across nearly all metrics and times, compared to the IFS ensemble baseline. The signature kernel score once again appears sharper, but has a notable shift to stronger predictions for later path lengths, as opposed to initial lead times, as seen clearly in Southern Hemisphere variables. Notably, the same model performance changes seen in geopotential in the Northern Hemisphere by CRPS and ENS MSE are also observed by the signature kernel. However, geopotential 850 in the Northern Hemisphere does not see a reversal in model performance by the signature kernel. This is due to the path length versus lead time comparison, where the change is small enough on the path length scale to not cause a complete reversal.
These behaviours reflect both the unique way the signature kernel score picks up on the structure of the data compared to existing metrics along with the full path length evaluation approach. These differences may support the use of the signature kernel in model evaluation alongside a standard lead time metric.
@@
Figure 2: Deterministic scorecard comparison of Fuxi and IFS Mean across weather variables and metrics, over three regions of the globe. Darker blues signify positive metric performance for Fuxi as compared to the IFS Mean metric value, evaluated across 15 days.
@@
Figure 3: Probabilistic scorecard comparison of Neural GCM and IFS ENS across weather variables and metrics, over three regions of the globe. Darker blues signify positive metric performance for Neural GCM as compared to the IFS ENS metric value, evaluated across 15 days.

### v2

@@ 2.3 Signature Augmentations
As previously mentioned, the current framing of the signature uniquely characterises a path up to translation and time parametrisation. Both of these pose issues for use as an evaluation metric and for use in a scoring rule. Learning the behaviour of weather path curvature is not enough if the forecast has significant bias via translation shifts. Weather has clear temporal cycles on the level of days, months, and years. The signature needs to relate path behaviour to time to evaluate a forecast. We introduce the following augmentations (Morrill et al., 2021). The basepoint augmentation $\phi^{b}:\mathcal{S}(\mathbb{R}^{d})\mapsto\mathcal{S}(\mathbb{R}^{d})$ is defined as:
@@
This augmentation adds another dimension to the path, which carries the time of the observation at that point. This allows for different or irregular sampling between forecast and observations. This is critical for weather forecasting, as forecasts and observations are regularly predicted and sampled at different intervals. Moreover, the effect of missing values is mitigated, as the time dimension aligns true positions with all existing values correctly. These augmentations combined remove any invariances and guarantee the uniqueness of the signature. This property is critical for the performance of a scoring rule involving the signature kernel, as it is required for strict propriety, discussed in the following subsection. See Appendix D for further discussion of signature augmentations.
@@ 2.4 The Signature Kernel Scoring Rule
The previous sections introduced the idea, intuition, and practical concerns with the signature of a path. Nonetheless, the signature remains an infinite sequence of real numbers, which still can’t be used in its current form. The primary concept involved is the kernel trick, which allows data to be analysed in a higher (or infinitely high) dimensional space, without explicitly computing the mapping, i.e. the full signature. A kernel is a positive semidefinite function that takes vectors in the original spaces as inputs, and returns the inner product of the vectors in the feature space (Wilimitis, 2019). We can define a mapping $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ to be a kernel if there exists a Hilbert space $\mathcal{H}$ (feature space), and a feature map $\phi:\mathcal{X}\rightarrow\mathcal{H}$, such that for all $x,z\in\mathcal{X}$ we have:
@@
which maps to a single score. However, to apply the kernel trick, the construction must take place in a reproducing kernel Hilbert space (RKHS). This guarantees the complete inner product structure that is required (Xu et al., 2023). If the original space $\mathcal{Y}$ is not itself a Hilbert space, we must first lift each path in $\mathcal{Y}$ to the RKHS $\mathcal{H}$ induced by a static kernel $k$ on $\mathcal{Y}$ via the canonical feature map. This step also allows us to incorporate a similarity structure on $\mathcal{Y}$. With sequential information, lifting the space before considering the signature would be a good learning strategy regardless of whether $\mathcal{Y}$ is already a Hilbert space (Salvi et al., 2021). Given an RKHS($\mathcal{H}$, $k$) with a kernel $k$ on $\mathcal{Y}$, the reproducing property lets us define a general static kernel as follows:
@@
In practice, this kernel is the solution of a Goursat partial differential equation (PDE) (Salvi et al., 2021). In fact, the previous condition of piecewise linearity was a necessary condition to be solved by a Goursat PDE. While other methods of solving the kernel exist, involving sequentialised kernels, the time complexity is at least $\mathcal{O}(dl^{2})$, for a path with dimension $d$ and length $l$ (Kiraly and Oberhauser, 2019). Provided the number of GPU threads exceeds the size of the discretisation grid chosen for PDE evaluation on (controlled by dyadic order), the time complexity reduces to $\mathcal{O}(dl)$ (Salvi et al., 2021). This breaks the quadratic relationship with path length, making the Goursat PDE approach the most suitable method for implementation, used by the associated sigkernel Python package. While the constant factor involved in solving the PDE means that evaluation takes longer in practice than point-wise metrics like RMSE or CRPS, the signature kernel scales at the same asymptotic rate. This property ensures it remains feasible for high-dimensional, long-range atmospheric data. See Appendix G for an additional discussion on discriminative power affecting training time. Finally, we can define the signature kernel score considering path data over the time interval $1:t$ and longitude and latitude $i,j$. Denoting the set of ensemble forecast paths as $F_{i,j,1:t}$ and the observation path as $o_{i,j,1:t}$, this score is as follows:
$\phi(F_{i,j,1:t},o_{i,j,1:t})=\mathbb{E}_{X,X^{\prime}\sim F_{i,j,1:t}}\left[K^{Sig}(X,X^{\prime})\right]-2\mathbb{E}_{X\sim F_{i,j,1:t}}\left[K^{Sig}(X,o_{i,j,1:t})\right]$ (1)
Using this score, we are considering both the spread of the forecasted paths in the first term and the closeness, measured by the signature kernel, of the forecasted paths to the true observed paths in the second term. We can note the similarity between this kernel score and the energy score, with the apparent sign switch as a result of kernels measuring similarity versus distance norms representing dissimilarity.
Theorem 1.
Under a characteristic static kernel, with paths of bounded variation on a compact domain, and uniqueness assured via augmentations, the signature kernel score described in Equation 1 is strictly proper.
The strict propriety of the score ensures truthful predictions, as the true path distribution is the unique minimiser of the expected score. Appendix E provides the full proof.
Additionally, we will define the kernel distance as the primary metric to be used on deterministic forecasts. Using the fact that the kernel defines an inner product distance on X and Y within the transformed space $\langle S(k_{x}),S(k_{y})\rangle$, we can instead consider the Euclidean distance on this space (Phillips and Venkatasubramanian, 2011). With a non-ensemble forecast $f_{i,j,1:t}$, the distance is without expectation, with the distance across time $1:t$ and longitude, latitude $i,j$ as:
@@
We now have a distance measure, as opposed to a kind of similarity measure calculated using the inner product of $\langle S(k_{x}),S(k_{y})\rangle$. Under this measure, all values are non-negative, with the distance equalling zero when observations and forecasts have the same path. Ignoring expectations, the difference between the score and distance is the addition of the term $K^{Sig}(Y,Y)$, which represents the self-similarity of the observation path. As will be shown, this will improve the interpretability of deterministic forecasts by generally making distance monotonic, with a greater focus on forecasting skill than forecast spread.
@@
for all $\mathbf{x},\mathbf{x^{\prime}}\in\mathbb{R}^{d}$. The RBF kernel changes how the data points in the sequence are related, by introducing a notion of locality, scaled by the hyperparameter $\sigma$. This changes how close values are in kernel space, which results in different behaviours being captured by the signature kernel. The RBF may be better suited to the non-linear relationships present in weather data, but requires greater consideration during implementation to correctly tune $\sigma$. When using the RBF, $\sigma$ was chosen to be equal to 1, following positive experimental results (Issa et al., 2023). Other kernels can be used, such as the more interpretable linear kernel, but it suffers significantly from numerical instability and requires much greater attention to the scale of the input data. Refer to Section 4.3 for further details.
Moreover, dyadic order for the PDE solver needs to be chosen. Higher dyadic orders result in higher accuracy, at the cost of calculation speed. Given our inputs are significantly scaled, such that a maximum value is roughly equal to 1, experimental results show that dyadic orders of 0 or 1 are likely sufficient (Salvi et al., 2021). We implement a dyadic order of 1, following previous work (Issa et al., 2023). The use of the Goursat PDE solver approach also affects what weather variables can be considered. In the setting of piecewise-linear fixed dimensional paths, which move in all dimensions, typically none of the elements of the signature are zero. This is referred to as a dense signature. On the other hand, rainfall data is generally zero-inflated, particularly for certain ranges of latitude-longitude coordinates. As a result, this may result in a sparse signature, with many terms equal to zero. Methods which handle sparse signatures require significant overhead, which can be avoided if we stay in contexts resulting in dense signatures (Reizenstein and Graham, 2018). Nevertheless, extending this work to explicitly accommodate sparsity remains a valuable direction for future work, as it would enable more robust handling of zero-inflated weather phenomena like precipitation.
@@
The signature kernel scoring rule and distance on probabilistic and deterministic forecasts respectively are implemented in a notably different way than standard scores. On WeatherBench 2, scores such as RMSE are calculated at every prediction time lag. Most models provide 30 to 60 values of forecasted data from their initialisation time, spaced by twelve or six hours apart, equivalent to 360 hours or 15 days. The maximal forecast lead time for a model corresponds to when performance is no longer significantly different from predicting the average climatology. These values are averaged across all initialisation times and all locations using latitude weighting, to report score against time lag. The ECMWF then calculates their headline scores by considering the maximal forecast lead time required to reach a certain score (ECMWF, 2022b).
@@
The signature kernel score can not be calculated this way. A path is a temporally related sequence, not a collection of values across space at one particular time. The signature kernel score is instead calculated across chronologically sequential paths of length $k$ of the predicted data. Instead of having a value calculated for each prediction time lag, a value is determined for each path length from the initialisation time. Similar to the previous scores, these values can be averaged across all initialisation times. Further details of implementation for model evaluation use can be found in Appendix F.
@@
The following two scorecards are created from a 10% initialisation sample evenly distributed across the year 2020. Scorecard variables are chosen based on the variable overlap between models of the ECMWF headline variables (ECMWF, 2024). In Figure 1, we compare the FuXi Weather model and IFS ENS Mean across deterministic metrics and signature kernel distance with 60 timesteps of six hours, equivalent to 15 days. In Figure 2, we compare Neural GCM and IFS ENS across probabilistic metrics and signature kernel score with 30 timesteps of twelve hours, equivalent to 15 days. In both cases, IFS is taken as the baseline, with normalised percentage differences in metrics against its comparison being shown in colour. Darker blues signify greater positive performance for the target model as opposed to the baseline, with the opposite true of darker reds. Following ECMWF standards, calculations are split by region, with the Northern Hemisphere containing the latitude slices from 20 to 90 degrees, the Tropics containing -20 to 20, and the Southern Hemisphere containing -90 to -20. There is no overlap between regions.
@@
In Figure 1, the domain for anomaly cross correlation (ACC) and MSE is the 60 future lead times, while the signature kernel scoring rule (SIGK) is displayed against path length between 2 and 59. Despite this difference, there is general agreement across metrics over model performance at different times. However, some notable behavioural differences are detected. The signature kernel distance measure has generally sharper discrimination for positive FuXi performance, seen clearly in the Southern Hemisphere evaluation. The signature kernel also changes model performance evaluations in certain areas, unanticipated by the standard metrics. Northern and Southern Hemisphere 10 meter wind speed improves for FuXi performance for longer path lengths with signature kernel evaluation, differing from worse performance at later time leads for ACC and MSE. This implies model predictions have resulted in more appropriate path behaviours, despite specific values being worse.
@@
In Figure 2, the domain for continuous ranked probability score (CRPS) and ensemble (ENS) MSE is the 30 future lead times, while SIGK is displayed against path length between 2 and 29. In this scorecard, Neural GCM demonstrates superior performance across nearly all metrics and times, compared to the IFS ensemble baseline. The signature kernel score once again appears sharper, but has a notable shift to stronger predictions for later path lengths, as opposed to initial lead times, as seen clearly in Southern Hemisphere variables. Notably, the same model performance changes seen in geopotential in the Northern Hemisphere by CRPS and ENS MSE are also observed by the signature kernel. However, geopotential at 850 hPA in the Northern Hemisphere does not see a reversal in model performance by the signature kernel. This is due to the path length versus lead time comparison, where the change is small enough on the path length scale to not cause a complete reversal. These behaviours reflect both the unique way the signature kernel score picks up on the structure of the data compared to existing metrics, along with the full path length evaluation approach. See Appendix G for an additional comparative simulation. These differences may support the use of the signature kernel in model evaluation alongside a standard lead time metric.
@@
Figure 1: Deterministic scorecard comparison of FuXi and IFS Mean across weather variables and metrics, over three regions of the globe. Darker blues signify positive metric performance for FuXi as compared to the IFS Mean metric value, evaluated over 15 days across geopotential (z) at 500 and 850 hPa, temperature (t) at 500 and 850 hPa, two meter temperature (2t), and 10 meter windspeed (10ws).
@@
Figure 2: Probabilistic scorecard comparison of Neural GCM and IFS ENS across weather variables and metrics, over three regions of the globe. Darker blues signify positive metric performance for Neural GCM as compared to the IFS ENS metric value, evaluated over 15 days across geopotential (z) at 500 and 850 hPa, temperature (t) at 500 and 850 hPa, two meter temperature (2t), and 10 meter windspeed (10ws).